OpenStudy (anonymous):

wy is an altitude of XYZ. What is the length of wx? the length of the altitude (wy) is 12. the length of wz is 22

OpenStudy (anonymous):

\[a^2=b^2+c^2\] so \[22^2=12^2+wx^2\]\[22^2-12^2=wx^2\]\[\sqrt{(22^2-12^2)} = wx =\sqrt{628}=25.06\]

OpenStudy (anonymous):

Here is a more accurate answer

OpenStudy (anonymous):

Hope that helps

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks